# Revealing Multimodal Contrastive Representation Learning through Latent Partial Causal Models

### v1

@@ A.4 The Proof of identifiability on hypersphere
We then connect the right-hand side of Eq. (18) with Eq. (30). To this end, since the two terms in the right-hand side of Eq. (18) are symmetrical, we focus on one of the two terms for convenience, e.g., $\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x})),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}$. Based on Lemma 4.1, it can be shown that:
$\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x})),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}$ (31)
$\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{t}\sim p(\mathbf{z}_{t}|\mathbf{z}_{x})}[-\log q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x})]\big{]}}$ (32)
$=$ $-\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}-\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau+\log C_{q}(\mathbf{z}_{x})\big{]}}$ (33)
$=$ $\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}-\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau\big{]}+\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x})\sim p(\mathbf{z}_{x})}[\log C_{q}(\mathbf{z}_{x})]}$ (34)
$=$ $\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}-\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau\big{]}+\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x})\sim p(\mathbf{z}_{x})}[\log\int e^{(\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau)}\mathrm{d}\mathbf{z}_{x}]}$ (35)
$=-\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}\big{(}\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\big{[}e^{\big{(}\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}$ (36)
@@
$\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{t})\sim p(\mathbf{z}_{t})}{\big{[}H(p(\mathbf{z}_{x}|\mathbf{z}_{t})),q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t}))\big{]}}=-\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}\big{(}\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\big{[}e^{\big{(}\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}.$ (37)
@@
A.4.2 Identifiability result on hypersphere
Theorem 3.1 represents a adaptation of Theorem 1 from Zimmermann et al. (2021) in the context of multi-modal setting. Specifically, within the confines of a single-modal framework, Theorem 3.1 is consistent with the findings presented in Theorem 1 in Zimmermann et al. (2021). Consequently, this alignment allows us to employ Propositions 1 and 2 from Zimmermann et al. (2021) to demonstrate that the global minimization of the objective outlined in Equation (6), as specified in Theorem 3.1, identifies the latent variables $\mathbf{z}_{x}$, as well as $\mathbf{z}_{x}$, up to linear transformations. For completeness, a brief proof is provided herein, with comprehensive details available in the original work. Clearly, the global minimum of the cross-entropy between two distributions is reached if they match by value and have the same support. Therefore, for the optimal solution of the objective loss Eq. (18) in Theorem 3.1, we have:
@@ A.4.2 Identifiability result on hypersphere
$C_{p}e^{(k\mathbf{z}^{T}_{x}\mathbf{z}_{x})}=C_{q}(\mathbf{z}_{x})e^{(\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{x})/\tau)},$
@@
here we can see that $\mathbf{h}$ maintains the dot product, which implies that $\mathbf{h}$ must be an orthogonal linear transformation by using Proposition 2 in Zimmermann et al. (2021). As a result, Theorem 3.1 is capable of identifying the latent variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$ up to an orthogonal linear transformation, i.e., the recovered latent variable $\mathbf{\hat{z}}_{x}$, obtained through the minimization of Equation (6), is linearly related to the true $\mathbf{z}_{x}$ as follows: $\mathbf{\hat{z}}_{x}=\mathbf{A}\mathbf{z}_{x}+\mathbf{c}$, where $\mathbf{A}$ represents an orthogonal matrix, and $\mathbf{c}$ is a constant vector.
A.5 The Proof of identifiability on convex bodies
A.5.1 The Proof of Theorem 3.2
@@ A.5 The Proof of identifiability on convex bodies
($\mathcal{L}$ converges to the symmetric cross-entropy) Under the assumptions defined in Eqs. (9)-(10) for the proposed latent partial causal model, the necessary condition $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, denoted as $\mathbf{h}$, for the optimal normalized symmetric contrastive loss given by Eq. (2) leads to the following reduction of the loss itself:
$\lim_{N\rightarrow\infty}\mathcal{L}-2\log N=\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x})),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}+\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{t})\sim p(\mathbf{z}_{t})}{\big{[}H(p(\mathbf{z}_{x}|\mathbf{z}_{t})),q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t}))\big{]}}$ (41)
@@
$q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t})=C_{q}(\mathbf{z}_{t})e^{-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau},$ (42)
$q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x})=C_{q}(\mathbf{z}_{x})e^{-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau},$ (43)
@@
Similar to the proof A.4.1, we first introduce the following Lemma.
@@
where $C^{\prime}=\int_{-\infty}^{\infty}\exp\big{(}-\delta(\mathbf{0},\mathbf{z}_{t})\big{)}\mathrm{d}\mathbf{z}_{t}$. It is obvious that (45) can be viewed as a Kernel Density Estimation over samples $\mathbf{z}_{x_{i}}\sim p(\mathbf{z}_{x})$, and obviously $\lim_{\tau\rightarrow 0_{+}}p(\mathbf{z}_{t})$ will converge to $p(\mathbf{z}_{x})$ (which is uniform distribution) under quite mild condition (for details of the convergence we refer to (Jiang, 2017)). ∎
The proof of Theorem 3.2 can be done by demonstrating that the right-hand side of Eq. (41) is equal to the right-hand side of Eq. (14) on convex bodies. To achieve this, we first demonstrate that the minimization of (14) implies that the following condition hold: $\mathbf{f}_{x}\circ\mathbf{g}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{f}_{t}\circ\mathbf{g}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})$, for all real pairs $\big{(}(\mathbf{m}_{x},\mathbf{z}_{x}),(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}$. In the second step, by the condition above, we show that $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, and they are independent of the modality-specific variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, respectively. Finally, by defining $\mathbf{h}=\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, and using the inference model (42) and (43), we obtain our result.
@@
On convex bodies, similar to Eqs. (21)-(26), and define $\mathbf{h}_{x}=\mathbf{f}_{x}\circ\mathbf{g}_{x}$ and $\mathbf{h}_{t}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, we can further rewrite the right-hand side of Eq. (14) as:
$-2\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{m}_{x},\mathbf{z}_{x},\mathbf{m}_{t},\mathbf{z}_{t})\sim p(\mathbf{m}_{x},\mathbf{z}_{x},\mathbf{m}_{t},\mathbf{z}_{t})}{\big{[}d\big{(}\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x}),\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}/\tau\big{]}}$ (46)
$-H(p(\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})))-D_{\text{KL}}(p(\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x}))||p(\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})))+\log Z_{\text{KDE}}$ (47)
$-H(p(\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})))-D_{\text{KL}}(p(\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t}))||p(\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})))+\log Z_{\text{KDE}}.$ (48)
Again, on convex bodies, since $d$ denote a distance metric, the first term (46) is minimized if and only if
[…]
In addition, the last two terms, (47) and (48) is minimized if both $\mathbf{h}_{x}$ and $\mathbf{h}_{t}$ map $(\mathbf{m}_{x},\mathbf{z}_{x})$ and $(\mathbf{m}_{t},\mathbf{z}_{t})$, respectively, to uniform variables on hypersphere (condition 2##\##, note that here uniform distributions also supports the symmetry of Eq. (42) and Eq. (43)). Since condition 1##\## does not affect the validity of condition 2##\##, it follows that condition 1##\## is a necessary condition for minimizing the right-hand side of Eq. (18).
@@
Similar to Step II in Appendix A.4.1, by differentiating the equation in the condition 1##\## with respect to $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, respectively, we can conclude that both $\mathbf{h}_{x}$ and $\mathbf{h}_{t}$ are independent of the modality-specific variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, respectively, i.e., $\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{h}_{x}(\mathbf{z}_{x})$ and $\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})=\mathbf{h}_{t}(\mathbf{z}_{t})$. Further, since $\mathbf{h}_{x}(\mathbf{z}_{x})=\mathbf{h}_{t}(\mathbf{z}_{t})$ hold, for all real pairs $(\mathbf{z}_{x},\mathbf{z}_{t})$ sampled from the conditional distribution $p(\mathbf{z}_{t}|\mathbf{z}_{x})$ defined in Eq. (9), this expression also holds true for $\mathbf{z}_{t}=\mathbf{z}_{x}$, which implies $\mathbf{h}_{x}(\mathbf{z}_{x})=\mathbf{h}_{t}(\mathbf{z}_{x})$. As a result, we can obtain: $\mathbf{h}_{x}=\mathbf{h}_{t}$.
Step III

### v2

@@ Appendix D The Proof of identifiability on hypersphere
We then connect the right-hand side of Eq. (14) with Eq. (20). To this end, since the two terms in the right-hand side of Eq. (14) are symmetrical, we focus on one of the two terms for convenience, e.g., $\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x})),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}$. Based on Lemma 4.1, it can be shown that:
$\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x})),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}$ (21)
$=$ $\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t}|\mathbf{z}_{x})}[-\log q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x})]\big{]}}$ (22)
$=$ $\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}-\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau+\log C_{q}(\mathbf{z}_{x})\big{]}}$ (23)
$=$ $\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}-\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau\big{]}+\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x})\sim p(\mathbf{z}_{x})}[\log C_{q}(\mathbf{z}_{x})]}$ (24)
$=$ $\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}-\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau\big{]}+\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x})\sim p(\mathbf{z}_{x})}[\log\int e^{(\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})/\tau)}\mathrm{d}\mathbf{z}_{x}]}$ (25)
Since $p(\mathbf{z}_{x})=|\mathcal{Z}|^{-1}$, and $p(\mathbf{z}_{t})=|\mathcal{Z}|^{-1}$ by Lemma 4.1, Eq. (25) simplifies to:
$=-\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}\big{(}\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\big{[}e^{\big{(}\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}+\log|\mathcal{Z}|$ (26)
On hypersphere space with radius $r$, due to $\lVert\mathbf{h}(\mathbf{z}_{x})-\mathbf{h}(\mathbf{z}_{t})\rVert=2r-2\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{t})$, Eq. 26 simplifies to:
$=\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\big{[}e^{-d\big{(}\mathbf{h}(\mathbf{z}_{x})\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}$ (27)
@@
$\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{t})\sim p(\mathbf{z}_{t})}{\big{[}H(p(\mathbf{z}_{x}|\mathbf{z}_{t})),q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t}))\big{]}}=$ $\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau\big{]}}$
$+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\big{[}e^{-d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}+\log|\mathcal{Z}|.$ (28)
@@
D.2 Identifiability result on hypersphere
Theorem 4.1 represents a adaptation of Theorem 1 from [67] in the context of multi-modal setting. Specifically, within the confines of a single-modal framework, Theorem 4.1 is consistent with the findings presented in Theorem 1 in [67]. Consequently, this alignment allows us to employ Propositions 1 and 2 from [67] to demonstrate that the global minimization of the objective outlined in Eq. (5), as specified in Theorem 4.1, identifies the latent variables $\mathbf{z}_{x}$, as well as $\mathbf{z}_{x}$, up to linear transformations. For completeness, a brief proof is provided herein, with comprehensive details available in the original work. Clearly, the global minimum of the cross-entropy between two distributions is reached if they match by value and have the same support. Therefore, for the optimal solution of the objective loss Eq. (14) in Theorem 4.1, we have:
@@ D.2 Identifiability result on hypersphere
$C_{p}^{-1}e^{(k\mathbf{z}^{T}_{x}\mathbf{z}_{x})}=C_{q}(\mathbf{z}_{x})^{-1}e^{(\mathbf{h}(\mathbf{z}_{x})^{T}\mathbf{h}(\mathbf{z}_{x})/\tau)},$
@@
here we can see that $\mathbf{h}$ maintains the dot product, which implies that $\mathbf{h}$ must be an orthogonal linear transformation by using Proposition 2 in [67]. As a result, Theorem 4.1 is capable of identifying the latent variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$ up to an orthogonal linear transformation, i.e., the recovered latent variable $\mathbf{f}_{x}(\mathbf{x})$, obtained through the minimization of Eq. (5), is linearly related to the true $\mathbf{z}_{x}$ as follows: $\mathbf{f}_{x}(\mathbf{x})=\mathbf{A}\mathbf{z}_{x}+\mathbf{c}$, where $\mathbf{A}$ represents an orthogonal matrix, and $\mathbf{c}$ is a constant vector.
Appendix E The Proof of identifiability on convex bodies
E.1 The Proof of Theorem 4.3
@@ Appendix E The Proof of identifiability on convex bodies
($\mathcal{L}$ converges to the symmetric cross-entropy) Under the assumptions defined in Eq. (7) for the proposed latent partial causal model, the necessary condition $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, denoted as $\mathbf{h}$, for the optimal normalized multimodal contrastiveloss given by Eq. (2) leads to the following reduction of the loss itself:
$\lim_{N\rightarrow\infty}\mathcal{L}-2\log N=\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x})),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}+\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{t})\sim p(\mathbf{z}_{t})}{\big{[}H(p(\mathbf{z}_{x}|\mathbf{z}_{t})),q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t}))\big{]}}$ (32)
@@
$q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t})=C_{q}(\mathbf{z}_{t})^{-1}e^{-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau},$ (33)
$q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x})=C_{q}(\mathbf{z}_{x})^{-1}e^{-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau},$ (34)
@@
Similar to the proof D.1, we first introduce the following Lemma.
@@
where $C^{\prime}=\int_{-\infty}^{\infty}\exp\big{(}-\delta(\mathbf{0},\mathbf{z}_{t})\big{)}\mathrm{d}\mathbf{z}_{t}$. It is obvious that (36) can be viewed as a Kernel Density Estimation over samples $\mathbf{z}_{x_{i}}\sim p(\mathbf{z}_{x})$, and obviously $\lim_{\tau\rightarrow 0_{+}}p(\mathbf{z}_{t})$ will converge to $p(\mathbf{z}_{x})$ (which is uniform distribution) under quite mild condition (for details of the convergence we refer to [30]). ∎
Proof sketch
Similar to hypersphere, the proof of Theorem 4.3 can be done by demonstrating that the right-hand side of Eq. (32) is equal to the right-hand side of Eq. (10) on convex bodies. To achieve this, using Lemma 5.1, we show that $\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, and they are independent of the modality-specific variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, respectively. Finally, by defining $\mathbf{h}=\mathbf{f}_{x}\circ\mathbf{g}_{x}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$, and using the inference model (33) and (34), we obtain our result.
@@
On convex bodies, and define $\mathbf{h}_{x}=\mathbf{f}_{x}\circ\mathbf{g}_{x}$ and $\mathbf{h}_{t}=\mathbf{f}_{t}\circ\mathbf{g}_{t}$. Consider C1 in Lemma 5.1, e.g., $\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})$ almost surely, for pair $\big{(}(\mathbf{m}_{x},\mathbf{z}_{x}),(\mathbf{m}_{t},\mathbf{z}_{t})\big{)}$. Similar to Step I in Appendix D.1, by differentiating it with respect to $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, respectively, we can conclude that both $\mathbf{h}_{x}$ and $\mathbf{h}_{t}$ are independent of the modality-specific variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, respectively, i.e., $\mathbf{h}_{x}(\mathbf{m}_{x},\mathbf{z}_{x})=\mathbf{h}_{x}(\mathbf{z}_{x})$ and $\mathbf{h}_{t}(\mathbf{m}_{t},\mathbf{z}_{t})=\mathbf{h}_{t}(\mathbf{z}_{t})$. Further, since $\mathbf{h}_{x}(\mathbf{z}_{x})=\mathbf{h}_{t}(\mathbf{z}_{t})$ hold, for all real pairs $(\mathbf{z}_{x},\mathbf{z}_{t})$ sampled from the conditional distribution $p(\mathbf{z}_{t}|\mathbf{z}_{x})$ defined in Eq. (7), this expression also holds true for $\mathbf{z}_{t}=\mathbf{z}_{x}$, which implies $\mathbf{h}_{x}(\mathbf{z}_{x})=\mathbf{h}_{t}(\mathbf{z}_{x})$. As a result, we can obtain: $\mathbf{h}_{x}=\mathbf{h}_{t}$.